# CFTs with Large Gap from Barnes-Wall Lattice Orbifolds

### v1

@@ 1.1. Overview
With this general summary out of the way, let us define the gap of a CFT. First take the case where $V$ is a holomorphic VOA. The gap of $V$ is then the smallest conformal weight $\mu_{V}$ of any non-vacuum Virasoro highest weight vector in $V$[Höh19]. Modularity of the partition function gives the bound $\mu_{V}\leq\lfloor\frac{c}{24}\rfloor+1$ [Höh96]. An extremal VOA (or CFT) is a holomorphic VOA that saturates that inequality [Höh96, Höh08, Wit07]. The most famous example of an extremal VOA is the Monster VOA $V^{\natural}$ [Bor86, FLM88]. It has central charge $c=24$ and $\mu_{V}=2$. Other extremal VOAs are known up to $c=40$, but no examples for higher central charge are known. In particular, no holomorphic VOA with $\mu_{V}>2$ is known.
Next consider non-holomorphic VOAs and their CFTs $\operatorname{{\mathcal{H}}}$. Here the gap is the smallest total conformal weight $h+\bar{h}$ of any non-vacuum vector in $\operatorname{{\mathcal{H}}}$ that is highest weight with respect to both the holomorphic and the anti-holomorphic Virasoro algebra $Vir\otimes\overline{Vir}$. In the special case when $V$ is holomorphic, $\operatorname{{\mathcal{H}}}=V\otimes\bar{V}$ and the gap is simply given by $\mu_{V}$ as above. So far, no full CFT with gap greater than 2 is known.
@@ 1.3. The steps in more detail
This article is structured in the following way. In section 2 we discuss some background on the construction of VOAs from lattices and orbifolds. In section 3 we introduce the Barnes-Wall lattices, their automorphism groups, the extraspecial 2-group $E(m)$ and its properties, and in particular its lift to automorphisms of the VOA. In section 4 we introduce the intertwining representation. We explain how to compute the intertwiner matrices explicitly and check that the 2-cocycle vanishes. In section 5 we then return to the construction of orbifold VOAs and apply these results to construct CFTs with large gaps.
@@ 2.1. Orbifold theory
where $\textrm{Irr}_{c_{g}}$ denotes the set of irreducible projective characters with $2$-cocycle $c_{g}$, $V_{[g,\chi]}$ denotes the irreducible projective representation of the centraliser $C_{G}(g)$ with character $\chi$ and $W_{[g,\chi]}$ is the corresponding irreducible module for the fixed-point subalgebra $V^{G}$. [DRX17, MT04] then establishes that all irreducible modules of $V^{G}$ are given by $W_{[g,\chi]}$ for some conjugacy class $g$ and some projective irreducible representation $\chi$ of $C_{G}(g)$. In summary, the irreducible modules are labeled by $[g,\chi]$ and can be obtained from the twisted modules using the projector
@@ 2.2. Lattice VOAs and twisted group algebras
is then an automorphism of the twisted lattice algebra, $\hat{g}\in Aut(\hat{L})$. Finally, $\hat{g}$ is called a standard lift if $u_{g}=1$ on the fixed point lattice $L_{g}$[Lep85].
@@ 2.3. Twisted modules and the defect representation
where $L^{\perp}_{g}$ is the lattice of points in $L$ orthogonal to the fixed point lattice $L_{g}$. $N$ is a finite Abelian group. To construct it explicitly, define $D:=1-g$ and find its Smith Normal Form
$S=PDQ\ ,\qquad S={\mathrm{diag}}(s_{1},s_{2},\ldots,s_{d},0,\ldots,0)\ ,$ (2.3.2)
with $s_{i}|s_{i+1}$ and $s_{i}\in\operatorname{\mathbb{Z}}_{+}$. If $\{\alpha_{1},\ldots,\alpha_{d},\delta_{1},\ldots,\delta_{\dim L-d}\}$ are the row vectors of $Q^{-1}$, then $N$ is generated by the $[\alpha_{i}]$ and has the form
@@
$C(\alpha,\beta)=\prod_{k=0}^{n-1}(-\xi_{n}^{k})^{-\langle\alpha g^{k}|\beta\rangle}$ (2.3.4)
@@ 3.1. Barnes-Wall lattices and their automorphism groups
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\qquad\sigma_{2}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\ .$ (3.1.1)
@@
$[z,{\Sigma_{ar}}]=1\qquad z^{2}={\Sigma_{ar}}^{2}=1\qquad[{\Sigma_{ar}},{\Sigma_{bs}}]=z^{\delta_{a,b}\delta_{r,s+1\mod 2}}\ ,$ (3.1.3)
where $[x,y]=xyx^{-1}y^{-1}$. The elements in $E(m)$ are then automorphisms of $BW(m)$. For the full automorphism group $BRW(m)$ we will follow [NRS01]. It can be obtained from the normalizer of $E(m)$ in $O(2^{m},\operatorname{\mathbb{R}})$. For $m>3$, its structure is
@@
More precisely, the real Clifford group $C_{m}$ is defined as the the normalizer of $E(m)$ in $O(2^{m},\operatorname{\mathbb{R}})$. To describe its structure, we first note that $E(m)/Z(E(m))\simeq\operatorname{\mathbb{F}}_{2}^{2m}$ is a vector space over $\operatorname{\mathbb{F}}_{2}$. This vector space has a quadratic form $Q$ defined as $Q(x)=x^{2}\in\operatorname{\mathbb{F}}_{2}$, where the square of $x\in E(m)$, which is ±$\pm\mathbb{1}$, is interpreted as ±$\pm 1\in\operatorname{\mathbb{F}}_{2}$. The automorphism group preserving $Q$ is $O^{+}(2m,2)$. The Clifford group is then given by $C_{m}=E(m).O^{+}(2m,2)$. By corollary 2.3 of [NRS01], it is generated by the following elements of $O(2^{m},\operatorname{\mathbb{R}})$:
@@ 3.2. Some properties of $E(m)$
the central product coming from the fact that $-\mathbb{1}$ is an element of both groups in the product. For the counting, we note that there are $2\cdot\binom{m}{k}3^{m-k}$ elements that have $k$ factors $\sigma_{1}\sigma_{2}$. Since an element of order 4 has to have an odd number of such factors, we have
@@
Define the Heisenberg representation (or Fock representation) ${F}$ of $G<O(V)$ as the graded representation
@@ 3.3. Lifting $E(m)$
So far we have only discussed automorphisms $g$ of the underlying Barnes-Wall lattices. To obtain automorphisms of the lattice VOA $V_{L}$, we first lift $g$ to an automorphism $\hat{g}$ of the twisted group algebra $\hat{L}$. Section 2.2 gives an explicit expression of how to do this by adjoining a function $u_{g}$. The lifted automorphism $\hat{g}$ is then also an automorphism of $V_{L}$; in fact, $Aut(V_{L})=N\cdot O(\hat{L})$ where $O(\hat{L})$ is the group of all such lifted lattice automorphism and $N$ is the inner automorphism group [DN99]. (In fact in our case $N<O(\hat{L})$ since our $L$ does not have vectors of length squared 2.)
@@
$[\hat{z},{\hat{\Sigma}_{ar}}]=1\qquad\hat{z}^{2}={\hat{\Sigma}_{ar}}^{2}=1\qquad[{\hat{\Sigma}_{ar}},{\hat{\Sigma}_{bs}}]=\hat{z}^{\delta_{a,b}\delta_{r,s+1\mod 2}}$ (3.3.3)
@@ 4.1. Intertwining representation and Schur averaging
To find $T$ we will use Schur averaging: Let $P$ and $P_{h}$ be the 2-cocycles of the representations $\rho$ and $\rho_{h}$. Since they are assumed to be projectively equivalent, we can find a function $B_{h}$ such that $PP_{h}^{-1}=dB_{h}$, giving the sought after coboundary. Let $t$ be a linear map. We can then take the Schur average of of $t$ as
@@ 4.2. Coordinate notation
and define the coordinate vector $a=({\underline{a}},{\underline{\check{a}}})$ with ${\underline{a}}=(a_{1},\ldots,a_{d})$ and ${\underline{\check{a}}}=(\check{a}_{1},\ldots,\check{a}_{d})\in\operatorname{\mathbb{F}}_{2}^{d}$. The idea is then to work in the exponent of $(-1)$ and reduce everything to a problem in linear algebra over $\operatorname{\mathbb{F}}_{2}$.
@@
$R_{1/2}=\begin{pmatrix}0&I\\ 0&0\end{pmatrix}\ .$ (4.2.3)
@@
$(\Delta^{h}_{1/2})_{ij}=\left\{\begin{array}[]{ccl}\Delta^{h}_{ij}&:&i<j\\ 0&:&\textrm{else}\end{array}\right.\ .$ (4.2.9)
@@ 4.3. Intertwining representation of $E(m)$
We now want to use the Schur average intertwiner operators (4.1.5) to construct a projective representation of $E(m)$ on $\Omega_{0}$. We proceed as follows: We will first construct intertwiner matrices $T_{h}$ for all $12m+$ 1 generators of $E(m)$. We then check that these indeed generate a projective representation by checking that they satisfy the projective version of the relations (3.3.3).
@@
$T_{ar}^{2}=\lambda(a,r)\mathbb{1}\ ,\qquad T_{ar}T_{bs}=\mu(a,r;b,s)T_{bs}T_{ar}$ (4.3.1)
@@
$a(\Delta^{h}_{1/2}+{\mathbb{h}}\Delta^{h}_{1/2}{\mathbb{h}}^{T})a^{T}=a(s_{h}^{T}+{\mathbb{h}}s_{h}^{T})$ (4.3.4)
@@
$a(\Delta^{h}_{1/2}+{\mathbb{h}}\Delta^{g}_{1/2}{\mathbb{h}}^{T}+\Delta^{g}_{1/2}+{\mathbb{g}}\Delta^{h}_{1/2}{\mathbb{g}}^{T})a^{T}=a(s_{h}^{T}+{\mathbb{h}}s_{g}^{T}+s_{g}^{T}+{\mathbb{g}}s_{h}^{T})$ (4.3.5)
@@ 5.2. The diagonal orbifold
We start from expression (2.1.4) for $\operatorname{{\mathcal{H}}}$. From the comments above we know that to this order only the untwisted characters contribute, and for those in turn only the Heisenberg character $\chi^{F}$ contributes. To establish the proposition, it is thus enough to compute how often the trivial representation appears in the appropriate products ${F}_{(i)}\otimes{F}_{(j)}$ of ${F}_{(0)},{F}_{(1)},{F}_{(2)}$ and ${F}_{(3)}$. The proposition then follows immediately from proposition 3.2. ∎

### v2

@@ 1.1. Overview
With this general summary out of the way, let us define the gap of a CFT. First take the case where $V$ is a holomorphic VOA. The gap of $V$ is then the smallest conformal weight $\mu_{V}$ of any non-vacuum Virasoro highest weight vector in VV[Höh19]. Modularity of the partition function gives the bound $\mu_{V}\leq\lfloor\frac{c}{24}\rfloor+1$ [Höh96]. An extremal VOA (or CFT) is a holomorphic VOA that saturates that inequality [Höh96, Höh08, Wit07]. The most famous example of an extremal VOA is the Monster VOA $V^{\natural}$ [Bor86, FLM88]. It has central charge $c=24$ and $\mu_{V}=2$. Other extremal VOAs are known up to $c=40$, but no examples for higher central charge are known. In particular, no holomorphic VOA with $\mu_{V}>2$ is known.
[…]
@@
Unfortunately this orbifold construction is rather involved. We therefore provide a less technical summary and outlook in the conclusion section at the end of the article. Readers who are not interested in the details of the construction of the CFT, but rather in its properties and possible applications, may want to skim through the rest of the introduction and then skip straight to section 6.
@@ 1.3. The steps in more detail
This article is structured in the following way. In section 2 we discuss some background on the construction of VOAs from lattices and orbifolds. In section 3 we introduce the Barnes-Wall lattices, their automorphism groups, the extraspecial 2-group $E(m)$ and its properties, and in particular its lift to automorphisms of the VOA. In section 4 we introduce the intertwining representation. We explain how to compute the intertwiner matrices explicitly and check that the 2-cocycle vanishes. In section 5 we then return to the construction of orbifold VOAs and apply these results to construct CFTs with large gaps. Finally in section 6 we give a less technical summary of the constructed CFT and its properties.
@@ 2.1. Orbifold theory
where $\textrm{Irr}_{c_{g}}$ denotes the set of irreducible projective characters with $2$-cocycle $c_{g}$, $V_{[g,\chi]}$ denotes the irreducible projective representation of the centraliser $C_{G}(g)$ with character $\chi$ and $W_{[g,\chi]}$ is the corresponding irreducible module for the fixed-point subalgebra $V^{G}$. [DRX17, MT04] then establishes that all irreducible modules of $V^{G}$ are given by $W_{[g,\chi]}$ for some conjugacy class $g$ and some projective irreducible representation $\chi$ of $C_{G}(g)$. In summary, the irreducible modules are labeled by $g,\chi]$ and can be obtained from the twisted modules using the projector
@@ 2.2. Lattice VOAs and twisted group algebras
[…]
@@ 2.3. Twisted modules and the defect representation
where $L^{\perp}_{g}$ is the lattice of points in $L$ orthogonal to the fixed point lattice $L_{g}$. $N$ is a finite Abelian group. To construct it explicitly, define D:=1gD:=1-g and find its Smith Normal Form
[…]
with $s_{i}|s_{i+1}$ and $s_{i}\in\operatorname{\mathbb{Z}}_{+}$. If $\{\alpha_{1},\ldots,\alpha_{d},\delta_{1},\ldots,\delta_{\dim L-d}\}$ are the row vectors of $Q^{-1}$, then $N$ is generated by the $\alpha_{i}]$ and has the form
@@
$C(\alpha,\beta)=\prod_{k=0}^{n-1}(-\xi_{n}^{k})^{-\langle\alpha g^{k}|\beta\rangle}\$(2.3.4)
@@ 3.1. Barnes-Wall lattices and their automorphism groups
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\qquad\sigma_{2}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\$. (3.1.1)
@@
$z,{\Sigma_{ar}}]=1\qquad z^{2}={\Sigma_{ar}}^{2}=1\qquad[{\Sigma_{ar}},{\Sigma_{bs}}]=z^{\delta_{a,b}\delta_{r,s+1\mod 2}}\ ,$ (3.1.3)
where $x,y]=xyx^{-1}y^{-1}$. The elements in $E(m)$ are then automorphisms of $BW(m)$. For the full automorphism group $BRW(m)$ we will follow [NRS01]. It can be obtained from the normalizer of $E(m)$ in $O(2^{m},\operatorname{\mathbb{R}})$. For $m>3$, its structure is
@@
More precisely, the real Clifford group $C_{m}$ is defined as the normalizer of $E(m)$ in $O(2^{m},\operatorname{\mathbb{R}})$. To describe its structure, we first note that $E(m)/Z(E(m))\simeq\operatorname{\mathbb{F}}_{2}^{2m}$ is a vector space over $\operatorname{\mathbb{F}}_{2}$. This vector space has a quadratic form $Q$ defined as $Q(x)=x^{2}\in\operatorname{\mathbb{F}}_{2}$, where the square of $x\in E(m)$, which is $\pm\mathbb{1}$, is interpreted as $\pm 1\in\operatorname{\mathbb{F}}_{2}$. The automorphism group preserving $Q$ is $O^{+}(2m,2)$. The Clifford group is then given by $C_{m}=E(m).O^{+}(2m,2)$. By corollary 2.3 of [NRS01], it is generated by the following elements of $O(2^{m},\operatorname{\mathbb{R}})$:
@@ 3.2. Some properties of $E(m)$
the central product coming from the fact that $\mathbb{1}$ is an element of both groups in the product. For the counting, we note that there are $2\cdot\binom{m}{k}3^{m-k}$ elements that have $k$ factors $\sigma_{1}\sigma_{2}$. Since an element of order 4 has to have an odd number of such factors, we have
@@
Define the Heisenberg representation (or Fock representation) F{F} of $G<O(V)$ as the graded representation
@@ 3.3. Lifting $E(m)$
[…]
@@
$\hat{z},{\hat{\Sigma}_{ar}}]=1\qquad\hat{z}^{2}={\hat{\Sigma}_{ar}}^{2}=1\qquad[{\hat{\Sigma}_{ar}},{\hat{\Sigma}_{bs}}]=\hat{z}^{\delta_{a,b}\delta_{r,s+1\mod 2}}$ (3.3.3)
@@ 4.1. Intertwining representation and Schur averaging
To find $T$ we will use Schur averaging: Let $P$ and $P_{h}$ be the 2-cocycles of the representations $\rho$ and $\rho_{h}$. Since they are assumed to be projectively equivalent, we can find a function $B_{h}$ such that $PP_{h}^{-1}=dB_{h}$, giving the sought after coboundary. Let $t$ be a linear map. We can then take the Schur average of $t$ as
@@ 4.2. Coordinate notation
and define the coordinate vector $a=({\underline{a}},{\underline{\check{a}}})$ with ${\underline{a}}=(a_{1},\ldots,a_{d})$ and ${\underline{\check{a}}}=(\check{a}_{1},\ldots,\check{a}_{d})\in\operatorname{\mathbb{F}}_{2}^{d}$. The idea is then to work in the exponent of (1)(-1) and reduce everything to a problem in linear algebra over $\operatorname{\mathbb{F}}_{2}$.
@@
$R_{1/2}=\begin{pmatrix}0&I\\ 0&0\end{pmatrix}\$. (4.2.3)
@@
$(\Delta^{h}_{1/2})_{ij}=\left\{\begin{array}[]{ccl}\Delta^{h}_{ij}&:&i<j\\ 0&:&\textrm{else}\end{array}\right.\$. (4.2.9)
@@ 4.3. Intertwining representation of $E(m)$
We now want to use the Schur average intertwiner operators (4.1.5) to construct a projective representation of $E(m)$ on $\Omega_{0}$. We proceed as follows: We will first construct intertwiner matrices $T_{h}$ for all $2m+1$ generators of $E(m)$. We then check that these indeed generate a projective representation by checking that they satisfy the projective version of the relations (3.3.3).
@@
$T_{ar}^{2}=\lambda(a,r)\mathbb{1}\ ,\qquad T_{ar}T_{bs}=\mu(a,r;b,s)T_{bs}T_{ar}\$(4.3.1)
@@
$a(\Delta^{h}_{1/2}+{\mathbb{h}}\Delta^{h}_{1/2}{\mathbb{h}}^{T})a^{T}=a(s_{h}^{T}+{\mathbb{h}}s_{h}^{T})\$(4.3.4)
@@
$a(\Delta^{h}_{1/2}+{\mathbb{h}}\Delta^{g}_{1/2}{\mathbb{h}}^{T}+\Delta^{g}_{1/2}+{\mathbb{g}}\Delta^{h}_{1/2}{\mathbb{g}}^{T})a^{T}=a(s_{h}^{T}+{\mathbb{h}}s_{g}^{T}+s_{g}^{T}+{\mathbb{g}}s_{h}^{T})\$(4.3.5)
@@ 5.2. The diagonal orbifold
[…]
@@ 5.3. The holomorphic orbifold
6. Conclusion
In this section we give a less technical summary of the conjectured CFT of gap 4, its construction and its known properties. We also speculate on some of its properties and discuss how one could establish them. The section is aimed at readers who are interested in using this CFT, but do not want to go through the details of its construction.
6.1. The construction
We constructed a holomorphic CFT by generalizing the construction of the Monster CFT. We started with a holomorphic lattice CFT obtained by compactifying 128 free bosons on the Barnes-Wall lattice $BW(7)$ of dimension 128. This lattice is even and self-dual, so that the resulting lattice CFT is indeed holomorphic. Moreover its shortest vectors have length squared 8, so that the lightest winding-momentum states of the CFT have weight 4.
Next we orbifolded the lattice CFT by the extraspecial group $E(7)$. This is a subgroup of the full automorphism group $BRW(7)$ of the lattice. It has $2^{15}$ elements. It is not Abelian, but it is as close to an Abelian group as possible in the sense that all its commutators commute with all group elements. As with all orbifolds, this leads to untwisted and twisted sectors. In the untwisted sector, the only states of weight less than 4 that survive the orbifold are the vacuum and its Virasoro descendants; all other states are projected out. For the twisted sectors, we established that all states have weight 4 or higher. In total our construction thus has no primary fields of weight less than 4, giving a gap of 4.
One motivation for constructing this CFT is to interpret it in the context of holography as being dual to quantum gravity on $AdS_{3}$. Because the central charge is relatively small, its dual gravity theory would be in the highly quantum regime. Since the lightest primary field has twice the weight of the energy-momentum tensor, the lightest matter field in the gravity theory would appear at twice the AdS scale. The AdS scale and the matter scale would thus be separated, even though that separation is of course not parametrically large. We note that the CFT is not an extremal CFT and that thus the lightest primary fields do not describe BTZ black holes, which would appear at weight 6. Of course, in using this language we extrapolate from the classical to the highly quantum regime; to properly use such holographic language we would need a large $c$ limit of a family of such CFTs, which we currently do not have. Nonetheless, the conjectured CFT provides the first example of a CFT where the mass scale is strictly bigger than the AdS scale.
In the interest of using the conjectured CFT for applications, let us now gather further information on its structure.
6.2. Spectrum
First, let us discuss its spectrum. The lightest states are given by $|0\rangle,L_{-2}|0\rangle,L_{-3}|0\rangle$. The lightest primaries appear at weight 4. On the one hand, primaries in the untwisted sector appear: First, there are 2797 states coming from the weight 4 free boson descendants in the Heisenberg module that are invariant under the action of $E(7)$. Second, there are states coming from the vectors of the BW lattice of length squared 8. Their number is given by the number of orbits of such vectors under $E(7)$, since each such orbit corresponds to the state given by the sum of lattice vector states, which remains invariant under $E(7)$. To count the orbits, we use the fact that $BRW(7)$ is the normalizer of $E(7)$, so that it simply permutes $E(7)$ orbits with each other. Moreover $BRW(7)$ acts transitively on the shortest vectors of length squared 8 (see Theorem 11.2 in [Gri05]). It follows that all such orbits have the same length. Picking any specific orbit, a straightforward calculation gives length 256. There are thus $1260230400/256=4922775$ weight 4 states coming from lattice vectors.
On the other hand, there are also weight 4 states coming from the twisted sectors associated to elements conjugate to $\Sigma_{11}$} in $E(7)$. (The lightest states in the twisted sectors for the group elements $\mathbb{1}$ and ${\Sigma_{11}}{\Sigma_{12}}$ have weight 6 and 8 respectively.) From (3.2.4) there are $4^{7}+2^{7}-2=16510$ such elements, each of which lives in a conjugacy class of size 2 with respect to $E(7)$, giving 8255 twisted sectors. Each sector has $\sqrt{N}=2^{2^{7-3}}=2^{16}$ ground states, where the finite group $N$ is described at the end of section 4.4. This gives a total of $540999680$ twisted states at weight 4. Since the intertwining representation in the $\Sigma_{11}$} sectors is trivial, all these states survive the orbifold projection.
In principle one can continue this analysis also for weight 5 states. Since the partition function is a modular function that is completely determined by the first six terms, knowing the spectrum up to weight 5 is thus enough to fix it completely. In this way one could thus obtain the full spectrum of the CFT. We will leave this for future work.
6.3. Correlation functions
Next, let us discuss how to compute correlation functions in the CFT. As usual, this reduces to the problem of determining the 3pt-functions, that is the structure constants of the holomorphic CFT. For the untwisted sector, this is in principle straightforward: Here, it is enough to restrict to the $E(7)$ invariant subsector of the original lattice CFT, whose structure constants are known and are essentially fixed by charge conservation. In particular the correlation functions only involving the lightest states, i.e. Virasoro descendants, will be the same as always.
Correlation functions involving twisted sector fields are more difficult. Here the missing ingredients are the fusion rules between the twisted sectors. One way to obtain them would be to compute the $S$-matrices of the twisted sector characters under modular transformations and then apply the Verlinde formula. Unfortunately obtaining exact expressions for those characters and then extracting the $S$-matrices is very hard, since among other things it also requires constructing the action of the orbifold group on the twisted sectors. In principle we could obtain this by the methods described in section 4. However, instead of just computing a single matrix element we would now have to construct the entire representation matrices, which is of course computationally much harder. We therefore leave this for future work. Perhaps less ambitiously, one could investigate the structure of the group multiplication of the conjugacy classes of the twisted sectors, requiring that the identity appear. This would provide at least partial selection rules for the structure constants.
6.4. Symmetries
Finally, let us say a few words about the symmetry of the orbifold CFT $V^{orb(E(7)}$. In general, obtaining the automorphisms of an orbifold CFT is a difficult question for various reasons, as we will describe below.
Let us first focus on $E(7)$. Clearly the untwisted sector $V^{E(7)}$ and the twisted sectors will each have $E(7)$ symmetry separately. Moreover $E(7)$ will permute the twisted sectors. However, in order for the orbifold theory $V^{orb(E(7))}$ to have $E(7)$ symmetry, this permutation action has to be compatible with the fusion rules of the twisted sector, which we did not compute.
The next question is if the orbifold CFT has more symmetry than that. The first observation is that the orbifold CFT has a finite symmetry group since it has no states of weight 1. Next, since the underlying lattice has the much larger automorphism group $BRW(7)$, it is natural to speculate that the orbifold CFT also has at least $BRW(7)$ symmetry. In fact, $BRW(7)$ is the normalizer of $E(7)$, which means that its action is compatible with the orbifold projection onto $E(7)$ invariant states. In the untwisted sector, $BRW(7)$ is thus an automorphism, and it also permutes the twisted modules. However, one would again have to check if it also compatible with the fusion rules.
However, the issue is that $BRW(7)$ is the automorphism group of the lattice, and not of the CFT: the symmetry group of the CFT is rather the lift $\hat{BRW(7)}$. Unlike for $E(7)$, we did not establish that the lifted group of $BRW(7)$ remains the same, and in fact we see no reason why this should be. But if $E(7)$ is not a normal subgroup of $\hat{BRW(7)}$, then the discussion in the above paragraph does not apply.
Finally, there is also the possibility of extra automorphisms, that is symmetries of the orbifold CFT that do not come from automorphisms of the original CFT. The most famous example of this is of course the Monster CFT, where the original Conway group symmetry gets enhanced to the Monster group. To identify such symmetries, a much more detailed analysis of the resulting orbifold CFT would be necessary. We again leave this for future work.